Assume that we have the following derivatives portfolio:

A long futures contract on stock A, with strike price 90 and a long put option contract on the same stock, with the same strike price. The option's premium is 6 and is payed today. The risk free rate of interest is 5%, and the time of expiration is 1 year.

What will be the present value profit for the above derivatives portfolio if the stock's spot price is 116 at the time of expiration.

Answer with two decimals, and use dot as a decimal separator.
Answer #1

When spot price is at $116 at the time of expiration, then Profit/loss from future contract= 116-90=26

Profit/Loss from put option=Option Payoff-Premium Paid= 0-6 (As the spot price is higher than exercise price, the option will expire with no payoff)

=-6

Hence, net profit from the portfolio after 1 year= 26-6=$20

Hence, Present value of profit=20/(1+risk free rate)=20/1.05=$19.04